### v1

# Bounds for the relative class number problem for function fields

Abstract.
We establish bounds on a finite separable extension of function fields in terms of the relative class number, thus reducing the problem of classifying extensions with a fixed relative class number to a finite computation. We also solve the relative class number two problem in all cases where the base field has constant field not equal to $\mathbb{F}_{\_}2$.
This paper is the result of a collaboration initiated during the workshop “Number Theory in the Americas 2” held at Casa Matemática Oaxaca in September 2024, supported in part by BIRS-CMO, the Journal of Number Theory, the Clay Mathematics Institute, and the IMU Commission for Developing Countries. Kedlaya was additionally supported by NSF (grants DMS-2053473, DMS-2401536) and UC San Diego (Warschawski professorship). Rama was additionally partially supported by ANII FCE_1_2023_1_176497.
1. Introduction
The relative class number one problem for number fields (more specifically, totally imaginary quadratic extensions of totally real number fields) was introduced by Stark [11] as a generalization of the classical Gauss class number one problem for imaginary quadratic number fields. While this problem has attracted much attention and seen much progress, it remains open even under assumption of the generalized Riemann hypothesis.
By contrast, the relative class number one problem for function fields was first considered by Leitzel–Madan [9] but is now fully resolved [5, 6, 7]. It is therefore natural to consider the problem of classifying finite separable extensions of function fields (of curves over finite fields) with relative class number equal to a fixed integer $m>1$. It is also natural to restrict to extensions which are either constant (i.e., induced by an extension of the constant field) or purely geometric (i.e., with no change in the constant field), as any extension can be interpreted as a constant extension followed by a purely geometric extension.
In this paper, we make two contributions to this problem. First, we reduce the relative class number $m$ problem for function fields, for any fixed positive integer $m$, to an explicit finite computation by giving bounds on the relevant parameters. Second, we solve the relative class number 2 problem for function fields in all cases where the base field has constant field not equal to $\mathbb{F}_{\_}2$.
Our arguments depend on a number of calculations made using SageMath and Magma. These are collected in a series of Jupyter notebooks available from our GitHub repository:
https://github.com/sarangop1728/twice-class-number.
1.1. Bounds for the class number problem
Our bounds for the relative class number $m$ problem for general $m$ take the following form.
Theorem 1.1.
Let $F^{\prime}/F$ be a finite separable extension of function fields of degree $d$ with relative class number $m$. Assume either that $F^{\prime}/F$ is constant and $d>1$, or that $F^{\prime}/F$ is purely geometric and $g^{\prime}>g$. Let $\mathbb{F}_{\_}q$ be the constant field of $F$. Let $g$ be the genus of $F$. Let $g^{\prime}$ be the genus of $F^{\prime}$.
  1. (a)
    We have $q\leq(\sqrt{m}+1)^{2}$.
  2. (b)
    If $F^{\prime}/F$ is constant, then $d$ and $g$ are bounded by functions of the form $a+b\log m$ where $a,b$ are some explicit constants.
  3. (c)
    If $F^{\prime}/F$ is purely geometric, then $g$ and $g^{\prime}$ are bounded by functions of the form $a+b\log m$ where $a,b$ are some explicit constants.
Proof.
For (a), see Lemma 3.2 in the purely geometric case and Lemma 3.6 in the constant case. For (b), see Lemma 3.6 for the bound on $d$ and for the bound on $g$ when $q>4$; Lemma 3.7 for the bound on $g$ when $q\in\{3,4\}$; and Lemma 3.8 for the bound on $g$ when $q=2$. For (c), see Lemma 3.1 for the bounds when $q>4$; Lemma 3.3 when $q\in\{3,4\}$; and Lemma 3.4 when $q=2$. ∎
As an aside, note that the bound on $q$ is best possible: it is equivalent to $m\geq(\sqrt{q}-1)^{2}$, and by Deuring’s theorem this ensures that there is an elliptic curve $E$ over $\mathbb{F}_{\_}q$ with #$\#E(\mathbb{F}_{\_}q)=m$. This gives rise to two different extensions with relative class number $m$: the function field of $E$ over a rational function field, and the quadratic constant extension of the function field of the quadratic twist of $E$.
The techniques used in the above arguments depend heavily on $q$, the size of the constant field of $F$. When $q>4$, the bounds are easily derived by applying the Weil bounds to the Prym variety of the extension. When $q\in\{3,4\}$, we must factor in the existence of an elliptic curve of order 1; for the complementary factor one gets a Weil-type exponential lower bound on the order of the Prym variety in terms of its dimension (see lemma 2.1 for the version used here), and then one must argue that the contribution of the elliptic curve forces the base field $F$ to have “too many” places of degree 1 compared to its genus. When $q=2$, there is no exponential lower bound available, but one can instead directly pass from the knowledge that $m$ is small to a lower bound on some weighted count of places of low degree.
For comparison, we draw attention to recent work of Kadets–Keliher [8]. Their aim is to solve the much harder problem of finding all extensions of function fields where the relative class group has a fixed exponent; if successful this would also give bounds on the relative class number problem, but at present their methods are not quite able to achieve this. However, these methods are different enough from ours that it is worth investigating whether one can combine ideas for better results.
1.2. The relative class number 2 problem
Our second contribution is to solve the relative class number 2 problem for function fields in all cases where the base field has constant field not equal to $\mathbb{F}_{\_}2$.
Theorem 1.2.
Let $F^{\prime}/F$ be a finite separable extension of function fields with relative class number $2$. Let $\mathbb{F}_{\_}q$ be the constant field of $F$ and assume that $q>2$. Let $g$ be the genus of $F$. Let $g^{\prime}$ be the genus of $F^{\prime}$.
  1. (a)
    In all cases we have $q\in\{3,4,5\}$.
  2. (b)
    If $F^{\prime}/F$ is a constant extension, then $[F^{\prime}:F]=2$ and
    $(q,g)\in\{(5,1),(4,1),(3,1),(3,2)\}.$
  3. (c)
    If $F^{\prime}/F$ is a purely geometric extension (that is, $\mathbb{F}_{\_}q$ is also the constant field of $F^{\prime}$) and $g=0$, then
    $(q,g^{\prime})\in\{(5,1),(4,1),(3,1),(3,2)\}.$
  4. (d)
    If $F^{\prime}/F$ is a purely geometric extension and $g=1$, then
    $(q,g^{\prime})\in\{(5,2),(4,2),(3,2),(3,3),(3,4)\}.$
  5. (e)
    If $F^{\prime}/F$ is a purely geometric extension and $g\geq 2$, then
    $(q,g,g^{\prime})\in\{(5,2,3),(4,2,3),(4,2,4),(4,3,5),(3,2,3),(3,2,4),(3,3,5),(3,4,7)\}.$
    Moreover, $[F^{\prime}:F]=2$ except in one case where $(q,g,g^{\prime})=(3,2,4)$ and $F^{\prime}/F$ is a noncyclic cubic extension.
See Proposition 4.1 for an overview of the result and for further navigation into the proof (including tables of examples).
As in [5, 6, 7], the argument here mostly involves work over $\mathbb{F}_{\_}3$ and $\mathbb{F}_{\_}4$, including a number of exhaustive searches for cyclic covers of particular curves (using explicit class field theory as implemented in Magma) and a few arguments to rule out most noncyclic covers of degree 3 or 4. One advantage over the case $m=1$ is that we can easily rule out cyclic covers of prime degree greater than 2 (Lemma 2.9.) One disadvantage is that we do have to account for one noncyclic cubic cover; fortunately, we end up with enough circumstantial information about the cover that it is relatively straightforward to prove its existence and even to identify it explicitly (Lemma 7.3).
The case of the relative class number 2 problem in which the base field has constant field $\mathbb{F}_{\_}2$ is expected to be substantially harder due to the large genera of curves involved (although again there is no need to worry about cyclic covers of odd prime degree). Recall that in [5] one finds an example of a quadratic extension of relative class number 1 where the base field has genus 7, and the arguments in [7] to rule out further such extensions require a restricted census over the $\mathbb{F}_{\_}2$-points of the moduli space of curves of genus 7.
Acknowledgments
Thanks to Carel Faber, Everett Howe, and David Roe for their help with generating data on Jacobians of curves and importing the data into LMFDB. Thanks also to Edgar Costa for arranging access to computing resources of the Simons Collaboration on Arithmetic Geometry, Number Theory, and Computation.
2. General facts
Given a finite separable extension of function fields $F^{\prime}/F$, we write $C,C^{\prime}$ for the curves corresponding to $F,F^{\prime}$; $q,q^{\prime}$ for the orders of the base fields of $C,C^{\prime}$; $g,g^{\prime}$ for the genera of $C,C^{\prime}$; and $h_{\_}F,h_{\_}{F^{\prime}}$ for the class numbers of $F,F^{\prime}$. We write $J=J(C)$ and $J^{\prime}=J(C^{\prime})$ for the Jacobians of $C,C^{\prime}$, so that #$\#J(\mathbb{F}_{\_}{q})=h_{\_}F$ and #$\#J^{\prime}(\mathbb{F}_{\_}{q^{\prime}})=h_{\_}{F^{\prime}}$.
We say that $F^{\prime}/F$ is constant if $F^{\prime}=F\cdot\mathbb{F}_{\_}{q^{\prime}}$, and purely geometric if $q^{\prime}=q$. If neither occurs, then $F^{\prime}$ is a purely geometric extension of $F\cdot\mathbb{F}_{\_}{q^{\prime}}$, which is a constant extension of $F$; it thus suffices to study relative class number problems in these cases.
Let $A$ denote the Prym variety associated to the covering $C^{\prime}\to C$, i.e., the kernel of the map
$\operatorname{Res}_{\_}{\mathbb{F}_{\_}{q^{\prime}}/\mathbb{F}_{\_}q}(J^{\prime})\to J,$
where $\operatorname{Res}$ denotes the Weil restriction. We then have $\dim(A)=g^{\prime}-g$ and #$\#A(\mathbb{F}_{\_}q)=h_{\_}{F^{\prime}/F}$.
For $\bullet$ a curve or abelian variety, let $T_{\_}{\bullet,q^{n}}$ be the trace of the $q^{n}$-power Frobenius on $\bullet$ (acting on the first étale cohomology with coefficients in $\mathbb{Q}_{\_}\ell$ for any prime $\ell$ which is nonzero in the base field). Let $P_{\_}\bullet(T)$ denote the characteristic polynomial of Frobenius, which we also call the Weil polynomial associated to $\bullet$.
2.1. Bounds on Prym varieties
The fundamental tool in studying the relative class number for function fields is the Weil bound for the Prym variety $A$, plus some refinements for small $q$.
2.1.1. $q>2$
Lemma 2.1 (Exponential lower bounds).
Let $B$ be a simple abelian variety defined over $\mathbb{F}_{\_}q$ for $q>2$. Assume that $B$ is not in the isogeny classes 1.3.ad, 1.4.ae. Then
(2.1) #$\#B(\mathbb{F}_{\_}q)\geq b_{\_}q^{\dim(B)},$
where $b_{\_}3=1.359$, $b_{\_}4=2$, and $b_{\_}q=(\sqrt{q}-1)^{2}$ for $q>4$.
Proof.
When $q>4$, the result follows from the Weil bound. When $q=3,4$, the result follows from [4, Theorem 3.2] with the given exceptions. ∎
For $q=3,4$, we will use the nature of the exceptional cases in Lemma 2.1 to establish that when $A$ has small order, its Frobenius traces must be large. To wit, the Prym variety decomposes as
$A\sim E^{\dim(A)-\dim(B)}\times B,$
where $E$ an elliptic curve of order 1 and $B$ is an abelian variety of order $m$ not containing $E$ as a factor. It follows from this isogeny decomposition that
(2.2) $T_{\_}{A,q}$ $=(\dim(A)-\dim(B))T_{\_}{E,q}+T_{\_}{B,q},$
$\geq(\dim(A)-\dim(B))q-2\dim(B)\sqrt{q},$
$\geq q\dim(A)-(q+2\sqrt{q})\dim(B),$
where the middle inequality follows from $T_{\_}{E,q}=q$ and the Weil bound on the trace of $B$. By Lemma 2.1 we have
(2.3) $\dim(B)\leq\frac{\log(\#A(\mathbb{F}_{\_}q))}{\log(b_{\_}q)},$
where $b_{\_}3=1.359$ and $b_{\_}4=2$.
2.1.2. $q=2$
This approach cannot be directly adapted to $q=2$ because there are infinitely many simple abelian varieties of order 1 over $\mathbb{F}_{\_}2$ [10], so there is no hope of working with them concretely. Instead, we give a lower bound for a suitable linear combination of traces of an abelian variety depending on its order.
Lemma 2.2.
Let $A$ be an abelian variety over $\mathbb{F}_{\_}2$. Then
(2.4) $T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0887\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}}$
$\geq 1.1691\cdot\dim(A)-1.9169\cdot\log\#A(\mathbb{F}_{\_}2).$
Proof.
Write
#$\#A(\mathbb{F}_{\_}2)=P_{\_}A(1)=\prod_{\_}{i=1}^{\dim(A)}(1-\alpha_{\_}i)(1-\overline{\alpha}_{\_}i)=\prod_{\_}{i=1}^{\dim(A)}(3-\beta_{\_}i),$
where the pairs $\alpha_{\_}i,\overline{\alpha_{\_}i}$ run over conjugate pairs of Frobenius eigenvalues of $A$ and $\beta_{\_}i:=\alpha_{\_}i+\overline{\alpha_{\_}i}\in[-2\sqrt{2},2\sqrt{2}]$. In this notation,
$T_{\_}{A,2}$ =_$=\sum_{\_}{i=1}^{\dim(A)}\beta_{\_}i,$
$T_{\_}{A,2^{2}}$ =_$=\sum_{\_}{i=1}^{\dim(A)}(\beta_{\_}i^{2}-4),$
$T_{\_}{A,2^{3}}$ =_$=\sum_{\_}{i=1}^{\dim(A)}(\beta_{\_}i^{3}-6\beta_{\_}i),$
$T_{\_}{A,2^{4}}$ =_$=\sum_{\_}{i=1}^{\dim(A)}(\beta_{\_}i^{4}-8\beta_{\_}i^{2}+8).$
The claim then follows from the fact that for $x\in[-2\sqrt{2},2\sqrt{2}]$,
$x+0.2456(x^{2}-4)+0.0887(x^{3}-6x)+0.0295(x^{4}-8x+8)\geq 1.1691-1.9169\log(3-x),$
as may be verified rigorously using interval arithmetic (the difference between the two sides is bounded below by 0.0004). See the file lemma2-2.ipynb. ∎
By comparison, when #$\#A(\mathbb{F}_{\_}2)=1$, [5, Lemma 5.6] yields
$T_{\_}{A,2}+0.2314\cdot T_{\_}{A,2^{2}}+0.093\cdot T_{\_}{A,2^{3}}+0.044\cdot T_{\_}{A,2^{4}}\geq 1.2766\cdot\dim(A);$
this is slightly better in practice than specializing Lemma 2.2 to the case where #$\#A(\mathbb{F}_{\_}2)=1$, but the proof technique does not extend to #$\#A(\mathbb{F}_{\_}2)>1$.
2.2. Bounds on point counts
We will play bounds off against upper bounds on rational points derived from nonnegativity constraints.
Lemma 2.3.
Let $C$ be a curve of genus $g$ over $\mathbb{F}_{\_}q$, for $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$. Then
(2.5) #$\#C(\mathbb{F}_{\_}q)\leq c_{\_}1g+c_{\_}0$
for the following values of the constants
$q$ $c_{\_}1$ $c_{\_}0$
$3$ $1.91$ $5.52$
$4$ $2.31$ $7.31$
Proof.
We adapt [5, Lemma 4.1] using the following values of $x_{\_}1$ and $x_{\_}2$.
  • ($q=3$)
    We use $x_{\_}1=0.932$, $x_{\_}2=0.0344$.
  • ($q=4)$
    We use $x_{\_}1=0.8$, $x_{\_}2=0.02$. ∎
For explicit computations, we supplement with some upper bounds for the number of points #$\#C(\mathbb{F}_{\_}q)$ for a curve $C$ of genus $g$, taken from manYPoints.
$g$ $q=3$ $q=4$ $q=3^{2}$ $q=4^{2}$
$1$ $7$ $9$ $16$ $25$
$2$ $8$ $10$ $20$ $33$
$3$ $10$ $14$ $28$ $38$
$4$ $12$ $15$ $30$ $45$
$5$ $13$ $17$ $35$ $53$
$6$ $14$ $20$ $38$ $65$
$7$ $16$ $21$ $43$ $69$
$8$ $18$ $23$ $46$ $75$
$9$ $19$ $26$ $50$ $81$
Table 1. Bound for #$\#C(\mathbb{F}_{\_}q)$ for a curve $C$ of genus $g$.
For $q=2$, in order to compare against Lemma 2.2, we use a bound including some point counts over extension fields.
Lemma 2.4.
Let $C$ be a curve of genus $g$ over $\mathbb{F}_{\_}2$. Let $a_{\_}r$ be the number of degree-$r$ places of $C$. Then
$a_{\_}1+0.3542\cdot 2a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4\leq 0.7365g+6.5441.$
Proof.
We use the parameters $x_{\_}2=0.8$ and $x_{\_}3=0.6$. See [5, Lemma 4.2]. The calculations can be found in the file linear-programming.ipybn. ∎
2.3. Constant case
For $F^{\prime}/F$ constant of degree $d$,
$P_{\_}{C^{\prime}}(T^{d})=P_{\_}C(T)P_{\_}C(\zeta_{\_}dT)\cdots P_{\_}C(\zeta_{\_}d^{d-1}T),$
where $\zeta_{\_}d$ is a primitive $d$-th root of unity. Consequently,
(2.6) $T_{\_}{A,q^{i}}=\begin{cases}-T_{\_}{C,q^{i}}&i\not\equiv 0\pmod{d},\\ (d-1)T_{\_}{C,q^{i}}&i\equiv 0\pmod{d}.\end{cases}$
When studying the relative class number $m$ problem for $m>1$, the possible degrees of constant extensions are heavily restricted. We first examine this in the prime case.
Lemma 2.5.
If $F^{\prime}/F$ is a constant extension of prime degree $\ell>2$ with $g_{\_}F>0$, then $h_{\_}{F^{\prime}/F}\equiv 0,1\pmod{\ell}$.
Proof.
When $F^{\prime}/F$ is a constant extension, we may obtain $J^{\prime}$ from $J$ by base extension from $\mathbb{F}_{\_}q$ to $\mathbb{F}_{\_}{q^{\ell}}$, and so $h_{\_}{F^{\prime}/F}=\#J(\mathbb{F}_{\_}{q^{\ell}})/\#J(\mathbb{F}_{\_}q)$. Now,
#$\#J(\mathbb{F}_{\_}{q^{\ell}})/\#J(\mathbb{F}_{\_}q)=P_{\_}{C^{\prime}}(1)/P_{\_}C(1)=P_{\_}C(\zeta)P_{\_}C(\zeta^{2})\cdots P_{\_}C(\zeta^{\ell-1})$
is a norm of the field $\mathbb{Q}(\zeta)$, where $\zeta$ is a primitive $\ell$-th root of unity. We conclude that $h_{\_}{F^{\prime}/F}\equiv 0,1\pmod{\ell}$. ∎
We then record some consequences for the case $m=2$.
Corollary 2.6.
Let $F^{\prime}/F$ be a constant extension with $g_{\_}F>0$, relative class number $2$, and prime degree $\ell=[F^{\prime}:F]$. Then $\ell=2$.
Corollary 2.7.
Let $F^{\prime}/F^{\prime\prime}/F$ be a tower of proper constant extensions, with corresponding fields of constants $\mathbb{F}_{\_}{q^{\prime}}/\mathbb{F}_{\_}{q^{\prime\prime}}/\mathbb{F}_{\_}q$. Assume also that $g_{\_}F>0$ and that $h_{\_}{F^{\prime}/F}=2$. Then,
  1. (a)
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=1$, then $q^{\prime\prime}\leq 4$.
  2. (b)
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=2$, then $q\leq 4$.
Proof.
[…]
  1. (a)
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=1$, then [5, Theorem 1.1] gives that $q^{\prime\prime}\leq 4$.
  2. (b)
    If $h_{\_}{F^{\prime}/F^{\prime\prime}}=2$, then $h_{\_}{F^{\prime\prime}/F}=1$ and [5, Theorem 1.1] gives that $q\leq 4$. ∎
2.4. Purely geometric case
When $F^{\prime}/F$ is purely geometric, we can rewrite Lemma 2.1 as an upper bound on the dimension in terms of $m=h_{\_}{F^{\prime}}/h_{\_}F$:
(2.7) $g^{\prime}-g=\dim(A)\leq\dfrac{\log(m)}{\log(b_{\_}q)}.$
A crucial observation that we will need is the following upper bound on the trace of the Prym variety [5, Equation (5.3)].
Lemma 2.8 (Prym trace less than rational points on curve).
Suppose that $F^{\prime}/F$ is purely geometric. Then for every $r\geq 1$,
(2.8) _$\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)T_{\_}{A,q^{n}}\leq ra_{\_}r,$
where $\mu$ is the Möbius function, and $a_{\_}r$ is the number of degree-$r$ places of $C$. In particular,
(2.9) $T_{\_}{A,q^{r}}\leq\#C(\mathbb{F}_{\_}{q^{r}})=\sum_{\_}{n\mid r}na_{\_}n.$
Proof.
From the short exact sequence that defines $A$, we have that $T_{\_}{C^{\prime},q^{r}}=T_{\_}{C,q^{r}}+T_{\_}{A,q^{r}}$. Therefore,
$0\leq\#C^{\prime}(\mathbb{F}_{\_}{q^{r}})$ $=q^{r}+1-T_{\_}{C^{\prime},q^{r}},$
$=q^{r}+1-T_{\_}{C,q^{r}}-T_{\_}{A,q^{r}},$
=#$=\#C(\mathbb{F}_{\_}{q^{r}})-T_{\_}{A,q^{r}}.\qed$
Similarly, _$\sum_{\_}{n\mid r}\mu\mathopen{}\mathclose{{}\left(\frac{r}{n}}\right)\#C^{\prime}(\mathbb{F}_{\_}{q^{n}})$ equals $r$ times the number of degree-$r$ places of $C^{\prime}$, and hence is also nonnegative.
When $g\geq 2$ and $d=2$, from [5, (7.7)] and Equation 2.9 we have
(2.10) $T_{\_}{A,q^{2i}}-t\leq\#C(\mathbb{F}_{\_}{q^{2i}})-2\#C(\mathbb{F}_{\_}{q^{i}}),$
where $t$ is the number of geometric ramification points. By Riemann–Hurwitz, $t\leq 2(g^{\prime}-(d-1)(g-1))$ if $q$ is odd and $t\leq g^{\prime}-(d-1)(g-1)$ if $q$ is even.
For cyclic extensions, we have an analogue of Lemma 2.5
Lemma 2.9.
Let $F^{\prime}/F$ be a purely geometric extension which is cyclic of prime degree $\ell>2$. Then $h_{\_}{F^{\prime}/F}\equiv 0,1\pmod{\ell}$.
Proof.
Let $\zeta$ be a primitive $\ell$-th root of unity. Form the étale cohomology groups $H^{1}(C),H^{1}(C^{\prime})$ with coefficients in $\mathbb{Q}_{\_}{\ell^{\prime}}$ for some prime $\ell^{\prime}$. The quotient $H^{1}(C^{\prime})/H^{1}(C)$, as a representation of the cyclic group $\mathrm{C}_{\_}\ell$, splits into eigenspaces for the $\ell-1$ nontrivial characters of $\mathrm{C}_{\_}\ell$. Since Frobenius commutes with the action of $\mathrm{C}_{\_}\ell$, it respects the eigenspace decomposition; choosing one eigenspace and taking the product of $1-\alpha$ as $\alpha$ runs over the Frobenius eigenvalues in that eigenspace yields an element of $\mathbb{Q}(\zeta)$ with norm $h_{\_}{F^{\prime}/F}$. This proves the claim. ∎
2.5. Noncyclic extensions of degree 3 and 4
Following [6], we enumerate the possible shapes of noncyclic, purely geometric extensions $F^{\prime}/F$ of function fields of degree 3 or 4. In the following discussion, let $F^{\prime\prime}/F$ be the Galois closure of $F^{\prime}/F$.
For $d=3$, if $F^{\prime}/F$ is not cyclic, then $F^{\prime\prime}/F$ has Galois group $S_{\_}3$ and so contains a quadratic subextension $F^{\prime\prime\prime}/F$ (the quadratic resolvent of the original cubic extension). This extension may be either constant or purely geometric. However, if there exists a degree-1 place of $F$ which does not lift to a degree-1 place of $F^{\prime}$ (which is automatic if #$\#C^{\prime}(\mathbb{F}_{\_}q)<\#C(\mathbb{F}_{\_}q)$), then as in [6, Lemma 4.5], we see that the quadratic resolvent must be purely geometric. In this case, $F^{\prime\prime\prime}/F$ corresponds to a covering $C^{\prime\prime\prime}\to C$ whose Prym $A^{\prime}$ has the property that
$J(C^{\prime\prime})\sim J\times A^{2}\times A^{\prime}.$
For $d=4$, one possibility is that $F^{\prime}/F$ admits an intermediate subfield (this includes the cyclic case). Excluding this option leaves three other possibilities. One is that $F^{\prime\prime}/F$ has Galois group $A_{\_}4$; by [6, Lemma 4.5], this case can be excluded by ensuring that some degree-1 place of $F$ lifts to a degree-4 place of $F^{\prime}$. In the other two cases, $F^{\prime\prime}/F$ has Galois group $S_{\_}4$ and again has a quadratic resolvent $F^{\prime\prime\prime}/F$ which can be either constant or purely geometric.
3. Bounds for the relative class number problem
In this section, we give bounds on a finite separable extension $F^{\prime}/F$ in terms of the relative class number $m:=h_{\_}{F^{\prime}/F}$, in the constant case and the purely geometric case.
3.1. Purely geometric extensions
When $F^{\prime}/F$ is purely geometric, the relevant statement is a bound on $g,g^{\prime},q$ in terms of $m$.
3.1.1. $q>4$
We have the following bounds on $g$ and $g^{\prime}$.
Lemma 3.1.
Suppose that $F^{\prime}/F$ is purely geometric, $g^{\prime}>g$, and $q>4$. Then
  1. (1)
    If $g\leq 1$, we have
    $g^{\prime}\leq 1+\dfrac{\log(m)}{2\log(\sqrt{q}-1)}\leq 1+\dfrac{\log(m)}{2\log(\sqrt{5}-1)}\leq 1+0.106\log(m).$
  2. (2)
    If $g>1$ and $[F^{\prime}:F]=d$, then
    $g$ $\leq 1+\dfrac{\log(m)}{2(d-1)\log(\sqrt{q}-1)}\leq 1+\dfrac{\log(m)}{2(d-1)\log(\sqrt{5}-1)}\leq 1+0.106\log(m),$
    $g^{\prime}$ $\leq 1+\dfrac{\log(m)}{\log(\sqrt{q}-1)}\leq 1+\dfrac{\log(m)}{\log(\sqrt{5}-1)}\leq 1+0.212\log(m).$
Proof.
From the Weil lower bound (Equation 2.7), we deduce that
$g^{\prime}-g=\dim(A)\leq\dfrac{\log(m)}{2\log(\sqrt{q}-1)}.$
When $g\leq 1$, we obtain the upper bound in (1). When $g>1$, Riemann–Hurwitz gives $g^{\prime}-g\geq(d-1)(g-1)$, from which we obtain (2). ∎
We bound $q$ in a similar way.
Lemma 3.2.
Suppose that $F^{\prime}/F$ is purely geometric and $g^{\prime}>g$. Then
$q\leq m+1+\mathopen{}\mathclose{{}\left\lfloor 2\sqrt{m}}\right\rfloor.$
Proof.
From Equation 2.7 we obtain that
$1\leq g^{\prime}-g\leq\dfrac{\log(m)}{2\log(\sqrt{q}-1)}.$
Solving this inequality for $q$ yields the result. ∎
3.1.2. $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$
Lemma 3.3.
Suppose that $F^{\prime}/F$ is purely geometric, $g^{\prime}>g$, and $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$.
  1. (1)
    If $g\leq 1$, we have
    $g^{\prime}$ $\leq 1+\frac{1}{q}\mathopen{}\mathclose{{}\left(c_{\_}1+c_{\_}0+(q+2\sqrt{q})\dfrac{\log(m)}{\log(b_{\_}q)}}\right),$
    $\leq\begin{cases}3.4767+7.0244\log(m)&q=3,\\ 3.405+2.8854\log(m)&q=4.\end{cases}$
  2. (2)
    If $g>1$ and $[F^{\prime}:F]=d$, then
    $g$ $\leq 1+\dfrac{1}{q(d-1)-c_{\_}1}\mathopen{}\mathclose{{}\left(c_{\_}1+c_{\_}0+(q+2\sqrt{q})\dfrac{\log(m)}{\log(b_{\_}q)}}\right),$
    $\leq\begin{cases}7.8166+19.333\log(m)&q=3,\\ 6.6924+6.8294\log(m)&q=4.\end{cases}$
    $g^{\prime}$ $\leq 1+\dfrac{2}{q-c_{\_}1}\mathopen{}\mathclose{{}\left(c_{\_}1+c_{\_}0+(q+2\sqrt{q})\dfrac{\log(m)}{\log(b_{\_}q)}}\right),$
    $\leq\begin{cases}14.6332+38.666\log(m)&q=3,\\ 12.3848+13.6588\log(m)&q=4.\end{cases}$
Proof.
From Equations 2.3, 2.3 and 2.2 we have that:
$c_{\_}1g+c_{\_}0\geq\#C(\mathbb{F}_{\_}q)\geq T_{\_}{A,q}$ $\geq q\dim(A)-(q+2\sqrt{q})\dim(B),$
$\geq q(g^{\prime}-g)-(q+2\sqrt{q})\dfrac{\log(m)}{\log(b_{\_}q)}.$
Solving for $g^{\prime}-g$ we obtain:
$g^{\prime}-g\leq\frac{1}{q}\mathopen{}\mathclose{{}\left(c_{\_}1g+c_{\_}0+(q+2\sqrt{q})\dfrac{\log(m)}{\log(b_{\_}q)}}\right).$
If $g\leq 1$, then $g^{\prime}-1\leq g^{\prime}-g$ and we obtain (1). If $g>1$, then Riemann–Roch gives $(d-1)(g-1)\leq(g^{\prime}-g)$ and we obtain (2); to minimize $g^{\prime}$ we first take $d=2$ and then $g^{\prime}=2g-1$. ∎
One can improve this bound by distinguishing the cases $d=2$ and $d>2$, using Equation 2.10 in the former case. We will use this strategy in the computations for $m=2$.
3.1.3. $q=2$
In this case, we cannot separate the Prym variety in any useful way, so we instead appeal to Lemma 2.2 and Lemma 2.4.
Lemma 3.4.
Suppose that $F^{\prime}/F$ is purely geometric, $g^{\prime}>g$, and $q=2$.
  1. (1)
    If $g\leq 1$, we have
    $g^{\prime}\leq 8.9408+1.6397\log(m).$
  2. (2)
    If $g>1$ and $[F^{\prime}:F]=d$, then
    $g$ $\leq 1+\frac{1}{d-1.8033}(8.9408+1.6397\log(m))\leq 45.454+8.3661\log(m),$
    $g^{\prime}$ $\leq 1+\frac{2}{2-1.8033}(8.9408+1.6397\log(m))\leq 89.908+8.3661\log(m).$
Proof.
By Equation 2.9,
$1.2751(0.7365g+6.5441)$ $\geq 1.2751(a_{\_}1+0.3542\cdot 2a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4),$
$\geq 1.2751\cdot a_{\_}1+0.2751\cdot 2a_{\_}2+0.0877\cdot 3a_{\_}3+0.0295\cdot 4a_{\_}4,$
=#$=\#C(\mathbb{F}_{\_}2)+0.2456\cdot\#C(\mathbb{F}_{\_}{2^{2}})+0.0877\cdot\#C(\mathbb{F}_{\_}{2^{3}})+0.0295\cdot\#C(\mathbb{F}_{\_}{2^{4}}),$
$\geq T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0877\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}},$
$\geq 1.1691(g^{\prime}-g)-1.9169\log(m).$
Solving for $g^{\prime}-g$ we obtain
$g^{\prime}-g\leq 0.8033g+7.1375+1.6397\log(m).$
If $g\leq 1$, then $g^{\prime}-1\leq g^{\prime}-g$ and we obtain (1). If $g>1$, then Riemann–Roch gives $(d-1)(g-1)\leq(g^{\prime}-g)$ and we obtain (2); to minimize $g^{\prime}$ we first take $d=2$ and then $g^{\prime}=2g-1$. ∎
Again, one can improve this bound by distinguishing the cases $d=2$ and $d>2$, using Equation 2.10 in the former case; for $m=1$ this is the content of [5, Lemma 10.1].
3.2. Constant extensions
When $F^{\prime}/F$ is constant, the relevant statement is a bound on $d,g,q$ in terms of $m$.
3.3. $q>4$
We apply the Weil bounds as follows.
Lemma 3.5.
Let $B$ be an abelian variety over $\mathbb{F}_{\_}q$, with $q>4$, and let $m$ be a positive integer. Suppose that #$\#B(\mathbb{F}_{\_}{q^{d}})=m\cdot\#B(\mathbb{F}_{\_}q)$ for some $d\geq 2$. Then
$\mathopen{}\mathclose{{}\left(\frac{q^{d/2}-1}{q^{1/2}+1}}\right)^{2\dim B}\leq m.$
Proof.
We have
$(q^{d/2}-1)^{2\dim(B)}$$\leq\#B(\mathbb{F}_{\_}{q^{d}}),$
$=m\cdot\#B(\mathbb{F}_{\_}q)\leq m(q^{1/2}+1)^{2\dim B},$
from which we deduce the desired bound. ∎
We now apply Lemma 3.5 in the case when $B=J$ is the Jacobian of a curve $C/\mathbb{F}_{\_}q$, corresponding to a constant extension of function fields $F^{\prime}/F$ with relative class number $m$, so that
(3.1) $\mathopen{}\mathclose{{}\left(\frac{q^{d/2}-1}{q^{1/2}+1}}\right)^{2g}\leq m.$
Using this, we are able to find explicit bounds on $d$, $q$, and $g$, solely in terms of $m$.
Lemma 3.6.
Let $F^{\prime}/F$ be a constant extension of function fields of degree $d>1$ and relative class number $m$. Then
(3.2) $d$ $\leq 2\log_{\_}2((\sqrt{2}+1)\sqrt{m}+1),$
(3.3) $q$ $\leq(\sqrt{m}+1)^{2}.$
In addition, if $q>4$, then
(3.4) $0<g\leq\frac{\log(m)}{2\log(\sqrt{5}-1)}.$
Proof.
To bound $d$, we apply Equation 3.1 and extremize by taking $q=2,g=1$ to obtain
$m\geq\mathopen{}\mathclose{{}\left(\frac{2^{d/2}-1}{2^{1/2}+1}}\right)^{2}.$
To bound $q$, we apply Equation 3.1 and extremize by taking $d=2,g=1$ to obtain
$m\geq\mathopen{}\mathclose{{}\left(\frac{q-1}{q^{1/2}+1}}\right)^{2}=(\sqrt{q}-1)^{2}.$
To bound $g$, we apply Equation 3.1 and extremize by taking $d=2,q=5$ to obtain
$m\geq(\sqrt{5}-1)^{2g}.\qed$
3.3.1. $q\in\mathopen{}\mathclose{{}\left\{3,4}\right\}$
For $q\in\{3,4\}$, we must distinguish between the cases $d>2$, for which Lemma 3.5 gives a nontrivial bound on $g$, and the case $d=2$, for which no such bound can exist because of the existence of elliptic curves with $m=1$. We handle these cases by adapting the argument from the purely geometric case, but now with the bounds on traces replaced with the explicit formula Equation 2.6.
Lemma 3.7.
Let $F^{\prime}/F$ be a constant extension of function fields of degree $d>1$ and relative class number $m$.
  1. (1)
    If $d>2$ and $q=3$, then
    $0<g\leq 1.17\log(m).$
  2. (2)
    If $d>2$ and $q=4$, then
    $0<g\leq 0.6\log(m).$
  3. (3)
    If $d=2$ and $q=3$, then
    $1.40<g\leq 19.34\log(m)+$ 1.4 .
  4. (4)
    If $d=2$ and $q=4$, then
    $1.370<g\leq 6.83\log(m)+$ 1.37 .
Proof.
For $d>2$, we again apply Equation 3.1 and extremize by taking $d=3$ to obtain
$m\geq\mathopen{}\mathclose{{}\left(\frac{q^{3/2}-1}{q^{1/2}+1}}\right)^{2g}.$
For $d=2$, we apply Lemma 2.3 and Equations 2.2 and 2.3 to write
$c_{\_}1g+c_{\_}0$$\geq\#C(\mathbb{F}_{\_}q),$
$=q+1-T_{\_}{C,q},$
$=q+1+T_{\_}{A,q},$
$\geq q+1+qg-(q+2\sqrt{q})\frac{\log(m)}{\log(b_{\_}q)},$
and hence
$g\leq\frac{1}{q-c_{\_}1}\mathopen{}\mathclose{{}\left((q+2\sqrt{q})\frac{\log(m)}{\log(b_{\_}q)}+c_{\_}0-q-1}\right).\qed$
3.3.2. $q=2$
For $q=2$, we must distinguish between the cases $d>3$, for which Lemma 3.5 gives a nontrivial bound on $g$, and the cases $d=2$ and $d=3$, where again we must argue in terms of bounds on point counts.
Lemma 3.8.
Let $F^{\prime}/F$ be a constant extension of function fields of degree $d>1$ and relative class number $m$.
  1. (1)
    If $d>3$ and $q=2$, then
    $0<g\leq 2.31\log(m).$
  2. (2)
    If $d=3$ and $q=2$, then
    $0<g\leq 2.8971+1.254\log(m).$
  3. (3)
    If $d=2$ and $q=2$, then
    $0<g\leq 5.7135+3.3642\log(m).$
Proof.
For $d>3$, we again apply Equation 3.1 and extremize by taking $d=4$ to obtain
$m\geq\mathopen{}\mathclose{{}\left(\frac{2^{4/2}-1}{2^{1/2}+1}}\right)^{2g}.$
For $d=3$, by Lemma 2.2, Lemma 2.4, and (2.6), we have
$3.2728\displaystyle 0.8085g+$ 3.2728 $\geq 1.0977(0.7365g+6.5441)-3.9107,$
$\geq 1.0977(a_{\_}1+0.3542\cdot 2a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4)-3.1329,$
$\geq 1.0977\cdot a_{\_}1+0.2751\cdot 2a_{\_}2-0.1774\cdot 3a_{\_}3+0.0295\cdot 4a_{\_}4-3.1329,$
=(#$=(\#C(\mathbb{F}_{\_}2)-2-1)+0.2456(\#C(\mathbb{F}_{\_}{2^{2}})-2^{2}-1)$
$\quad+0.1774(2^{3}+1-\#C(\mathbb{F}_{\_}{2^{3}}))+0.0295(\#C(\mathbb{F}_{\_}{2^{4}})-2^{4}-1),$
$=-T_{\_}{C,2}-0.2456\cdot T_{\_}{C,2^{2}}+0.0887\cdot 2T_{\_}{C,2^{3}}-0.0295\cdot T_{\_}{C,2^{4}},$
$=T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0887\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}},$
$\geq 1.1691\dim(A)-1.9169\log\#A(\mathbb{F}_{\_}2),$
$=1.1691\cdot 2g-1.9169\log(m).$
For $d=2$, we have similarly
$3.2555\displaystyle 0.5993g+$ 3.2555 $\geq 0.8136(0.7365g+6.5441)-2.0688,$
$\geq 0.8136(a_{\_}1+0.3542\cdot a_{\_}2+0.1524\cdot 3a_{\_}3+0.0677\cdot 4a_{\_}4)-2.0688,$
$=0.8136\cdot a_{\_}1-0.2751\cdot 2a_{\_}2+0.0887\cdot 3a_{\_}3-0.0295\cdot 4a_{\_}4-2.0688,$
=(#$=(\#C(\mathbb{F}_{\_}2)-2-1)+0.2456(2^{2}+1-\#C(\mathbb{F}_{\_}{2^{2}}))$
$\qquad+0.0887(\#C(\mathbb{F}_{\_}{2^{3}})-2^{3}-1)+0.0295(2^{4}+1-\#C(\mathbb{F}_{\_}{2^{4}})),$
$=-T_{\_}{C,2}+0.2456\cdot T_{\_}{C,2^{2}}-0.0887\cdot T_{\_}{C,2^{3}}+0.0295\cdot T_{\_}{C,2^{4}},$
$=T_{\_}{A,2}+0.2456\cdot T_{\_}{A,2^{2}}+0.0887\cdot T_{\_}{A,2^{3}}+0.0295\cdot T_{\_}{A,2^{4}},$
$\geq 1.1691\dim(A)-1.9169\log\#A(\mathbb{F}_{\_}2),$
$=1.1691g-1.9169\log(m).\qed$
4. $m=2$: an overview
We next turn to the case $m=2$. In the next three sections, we give a proof of the following result summarized in Table 2.
Proposition 4.1.
Let $F^{\prime}/F$ be a finite separable extension of function fields with $h_{\_}{F^{\prime}}/h_{\_}F=2$, which is either constant or purely geometric; assume in addition that $q>2$.
  • We have $q\leq 5$.
  • If $F^{\prime}/F$ is constant, then $d=2$ and the isogeny class of $J$ is one of 1.5.e (for $q=5$), 1.4.d (for $q=4$), or 1.3.d or 2.3.ei (for $q=3$).
  • If $F^{\prime}/F$ is purely geometric and $g=0$, then the isogeny class of $J^{\prime}$ is one of 1.5.ae (for $q=5$), 1.4.ad (for $q=4$), or 1.3.ad or 2.3.aei (for $q=3$).
  • If $F^{\prime}/F$ is purely geometric and $g=1$, then the isogeny classes of $J^{\prime}$ and $C$ form a pair listed in Table 3.
  • If $F^{\prime}/F$ is purely geometric, $g\geq 2$, and $d=2$, then $(q,g,g^{\prime})$ must equal a tuple indicated in Table 2.
  • If $F^{\prime}/F$ is purely geometric, $g\geq 2$, and $d>2$, then $F^{\prime}/F$ is equal to a certain noncyclic extension with $q=3,d=3,g=2,g^{\prime}=4$; see Lemma 7.3.
Proof.
The case $q>4$ is treated in section 5; the case $q=4$ is treated in section 6; the case $q=3$ is treated in section 7. More precise cross-references can be found in Table 2. ∎
Proposition 4.1 reduces the relative class number 2 problem for function fields with $q>2$ to an exhaust over cyclic covers for some particular values of $q,g,d,g,g^{\prime}$ with $g\geq 2$. We discuss this exhaust in section 8.
Case $g$ $g^{\prime}$ Proof
$q>4$ constant $1$ 5.2
purely geometric $2$ $3$ 5.3
$q=4$ constant $1$ 6.2
$g\leq 1$ $1$ $2$ 6.3.1, 6.3.2
$d=2$ $2$ 33334444 6.3.3
$3$ $5$
$d>2$ 6.3.4
$q=3$ constant $2$ 7.2
$g\leq 1$ $1$ $4$ 7.3.1, 7.3.2
$d=2$ $2$ 33334444 7.3.3
$3$ 55556666
$4$ $7$
$d=3$ $2$ $4$ 7.3.4
$d>3$ 7.3.5
Table 2. Bounds for the relative class number 2 problem.
$q$ Isogeny class of $J^{\prime}$ Isogeny class of $C$
5 2.5.ag_s 1.5.ac
2.5.af_o 1.5.ab
2.5.ae_k 1.5.a
2.5.ad_g 1.5.b
2.5.ac_c 1.5.c
2.5.ab_ac 1.5.d
2.5.a_ag 1.5.e
4 2.4.ae_l 1.4.ab
2.4.ad_i 1.4.a
2.4.ac_f 1.4.b
2.4.ab_c 1.4.c
2.4.a_ab 1.4.d
3 2.3.aek 1.3.ac
2.3.acg 1.3.a
2.3.abe 1.3.b
2.3.ac 1.3.c
2.3.ba 1.3.d
3.3.aekas 1.3.b
3.3.adfag 1.3.c
3.3.acag 1.3.d
3.3.aelay 1.3.a
3.3.adhaq 1.3.b
3.3.acdai 1.3.c
3.3.ababa 1.3.d
4.3.aefaag 1.3.d
Table 3. Pairs of isogeny classes for $(J^{\prime},C)$ corresponding to maps $C^{\prime}\to C$ of curves over $\mathbb{F}_{\_}q$ of relative class number two, when $C$ is an elliptic curve.
5. $m=2$ and $q>4$
We prove Proposition 4.1 for $q>4$.
5.1. Analysis of the Prym
From the Weil bound, if #$\#A(\mathbb{F}_{\_}q)=2$, then $q=5$ and $\dim(A)=1$. There is a unique isogeny class satisfying these requirements, namely 1.5.ae.
5.2. Constant case
In this case, by Corollary 2.7, $d$ must be prime; by Corollary 2.6, we must in fact have $d=2$. Consequently, $A$ must be the quadratic twist of $J=C$. This implies that $P_{\_}C(T)=T^{2}+2T+5$, and $C$ is isogenous to the elliptic curve with label 1.5.e.
5.3. Purely geometric case
In this case, Riemann-Hurwitz gives the inequality
$(d-1)(g-1)\leq 1,$
implying that either $g\leq 1$ or $g=d=2$. We separate into cases based on $g$.
5.3.1. $g=0$
In this case, we have that the Prym variety $A$ is isogenous to $J^{\prime}$. Since $h_{\_}{F^{\prime}/F}=h_{\_}{F^{\prime}}=\#A(\mathbb{F}_{\_}5)=\#J^{\prime}(\mathbb{F}_{\_}5)$, the analysis of Section 5.2 gives that $J^{\prime}=C^{\prime}$ must be an elliptic curve in the isogeny class 1.5.ae.
5.3.2. $g=1$
In this case, $J^{\prime}\sim J\times A=C\times A$. In particular, $A$ is an elliptic curve in the isogeny class 1.5.ae, and for every Jacobian of the form $J^{\prime}=C\times A$, we obtain a covering $C^{\prime}\to C$ by composing the Abel–Jacobi map with the natural projection. To find these coverings, we loop through all isogeny classes of elliptic curves $C$ over $\mathbb{F}_{\_}5$ and save the ones for which $C\times A$ is a Jacobian. This LMFDB query confirms the results of Table 3.
5.3.3. $g=2$
In this case, $d=2$, $g^{\prime}=3$, and there is nothing more to check now.
6. $m=2$ and $q=4$
We prove Proposition 4.1 for $q=4$.
6.1. Analysis of the Prym
Up to isogeny, we can write
$A\sim E^{\dim(A)-\dim(B)}\times B,$
where $E$ is the elliptic curve with #$\#E(\mathbb{F}_{\_}4)=1$ (isogeny class 1.4.ae) and $B$ is an abelian variety not containing $E$ as a factor and satisfying #$\#B(\mathbb{F}_{\_}4)=2$. By [4, Theorem 3.2], we have #$\#B(\mathbb{F}_{\_}4)\geq 2^{\dim(B)}$ and so $\dim(B)=1$; specifically, $B$ is an elliptic curve with #$\#B(\mathbb{F}_{\_}4)=2$ (isogeny class 1.4.ad). We have
$T_{\_}{E,4}=4,\qquad T_{\_}{B,4}=3,$
$T_{\_}{E,4^{2}}=8,\qquad T_{\_}{B,4^{2}}=1,$
and so
(6.1) $T_{\_}{A,4}=4\dim(A)-1,\qquad T_{\_}{A,4^{2}}=8\dim(A)-7.$
6.2. Constant case
By Corollary 2.7, the only way for $F^{\prime}/F$ to be constant of composite degree would be to have an intermediate field $F^{\prime\prime}/F$ with $[F^{\prime\prime}/F]=1$; but in this case, we would have $q^{\prime\prime}\geq 4^{2}>5$. Consequently, if $F^{\prime}/F$ is constant, then $d=2$.
We have $T_{\_}{A,4}=-T_{\_}{C,4}=\#C(\mathbb{F}_{\_}4)-q-1$ with $\dim(A)=g$. By Equation 6.1,
#$\#C(\mathbb{F}_{\_}4)=5+4\dim(A)-1=4+4g.$
Using Lemma 2.3 we get $7.314+4g=\#C(\mathbb{F}_{\_}4)\leq 2.31g+$ 7.31 from which we can bound $g\leq 1$, so $g=1$. Since $A$ is the quadratic twist of $C$, this implies that $P_{\_}C(T)=T^{2}+3T+4$, and $C$ is isogenous to the elliptic curve with label 1.4.d.
6.3. Purely geometric case
By Equation 2.9, Equation 6.1, and Lemma 2.3,
(6.2) $7.314(g^{\prime}-g-1)+3=T_{\_}{A,4}\leq\#C(\mathbb{F}_{\_}4)\leq 2.31g+$ 7.31 .
6.3.1. $g=0$
In this case, Equation 6.2 gives $4g^{\prime}-1\leq 5=\#C(\mathbb{F}_{\_}4)$, so that $g^{\prime}=1$. Therefore, $C^{\prime}$ is isogenous to $B$ itself.
6.3.2. $g=1$
When $g=1$, we get $59\geq 4g^{\prime}-$ 5 and so $g^{\prime}\leq 3$.
For $g^{\prime}=2$, we have $J^{\prime}\sim J\times B$; according to this LMFDB query, the possible isogeny classes for $J^{\prime}$ are 2.4.ae_l, 2.4.ad_i, 2.4.ac_f, 2.4.ab_c, 2.4.a_ab.
For $g^{\prime}=3$, we have $J^{\prime}\sim J\times E\times B$. This LMFDB query shows that there are no Jacobians of this form.
This confirms the values listed in Table 3.
6.3.3. $g\geq 2$ and $d=2$
In this case,
$2.31g+7.31\geq\#C(\mathbb{F}_{\_}4)\geq T_{\_}{A,4}\geq 4(g-1)-1=4g-5,$
and so $g\leq 7$. In addition, by Lemma 2.3 and Equation 2.10,
#$\#C(\mathbb{F}_{\_}{4^{2}})$ $\geq 2\#C(\mathbb{F}_{\_}4)+T_{\_}{A,4^{2}}-t\geq 2\#C(\mathbb{F}_{\_}4)+8\dim(A)-7-t\geq 2\#C(\mathbb{F}_{\_}4)+8g-15,$
$\geq 2T_{\_}{A,4}+8g-15=8\dim(A)-2+8g-15\geq 16g-25.$
From Table 1, we see that for $g=7$, #$\#C(\mathbb{F}_{\_}{4^{2}})\leq 69<87=16g-25$; for $g=6$, #$\#C(\mathbb{F}_{\_}{4^{2}})\leq 65<71=16g-25$; and for $g=5$, #$\#C(\mathbb{F}_{\_}{4^{2}})\leq 53<55=16g-25$.
For $g=4$, we have
$45\geq\#C(\mathbb{F}_{\_}{4^{2}})\geq 16g-25+15(g^{\prime}-2g+1)=39+15(g^{\prime}-2g+1),$
and so $g^{\prime}=7$ and $F^{\prime}/F$ is unramified. We may rule out this case as follows. Since $F^{\prime}/F$ is unramified, #$\#C^{\prime}(\mathbb{F}_{\_}4)\equiv 0\pmod{2}$ and so #$\#C(\mathbb{F}_{\_}4)-T_{\_}{A,4}=\#C(\mathbb{F}_{\_}4)-11$ is a nonnegative even integer. Since $1745\geq\#C(\mathbb{F}_{\_}{4^{2}})\geq 2\#C(\mathbb{F}_{\_}4)+$ 17 by Table 1 and #$\#C(\mathbb{F}_{\_}{4^{2}})\equiv\#C(\mathbb{F}_{\_}4)\pmod{2}$, the pair (#$(\#C(\mathbb{F}_{\_}4),\#C(\mathbb{F}_{\_}{4^{2}}))$ must be one of
[…]
The indicated LMFDB searches show that none of these cases can occur.
For $g=3$, we have $114\geq\#C(\mathbb{F}_{\_}4)\geq 4(g^{\prime}-g)-$ 1. For the exhaust, we record the constraint
(6.3) $(q,g,g^{\prime})=(4,3,5)\Longrightarrow\#C(\mathbb{F}_{\_}4)\geq 7,\#C(\mathbb{F}_{\_}{4^{2}})\geq 2\#C(\mathbb{F}_{\_}4)+9.$
If $g^{\prime}=6$, this improves to #$\#C(\mathbb{F}_{\_}4)\geq 11$ and #$\#C(\mathbb{F}_{\_}{4^{2}})\geq 38$. Since $38\geq\#C(\mathbb{F}_{\_}{4^{2}})$ by Table 1 and #$\#C(\mathbb{F}_{\_}{4^{2}})\equiv\#C(\mathbb{F}_{\_}4)\pmod{2}$, the pair (#$(\#C(\mathbb{F}_{\_}4),\#C(\mathbb{F}_{\_}{4^{2}}))$ must be one of
[…]
The indicated LMFDB searches show that none of these cases can occur.
For $g=2$, we have $110\geq\#C(\mathbb{F}_{\_}4)\geq 4(g^{\prime}-g)-$ 1 and so $g^{\prime}\leq 4$. For the exhaust, we record the constraint
(6.4) $(q,g,g^{\prime})=(4,2,4)\Longrightarrow\#C(\mathbb{F}_{\_}4)\geq 7.$
6.3.4. $g\geq 2$ and $d>2$
In this case, by Riemann–Hurwitz, Equation 6.1, and Lemma 2.3,
$2.31g+7.31\geq\#C(\mathbb{F}_{\_}4)\geq T_{\_}{A,4}\geq 4(d-1)(g-1)-1\geq 8(g-1)-1,$
which implies $g=2$. Now from Table 1 we have
$10\geq\#C(\mathbb{F}_{\_}4)\geq T_{\_}{A,4}=4\dim(A)-1,$
and so $\dim(A)\leq 2$; consequently, $d=3$ and the extension is unramified.
By Lemma 2.9, $F^{\prime}/F$ cannot be cyclic. As in section 2.5, the quadratic resolvent of the Galois closure $F^{\prime\prime}/F$ must be purely geometric. The Prym of the quadratic resolvent is an elliptic curve $E^{\prime}$ such that
$0\leq\#C^{\prime\prime}(\mathbb{F}_{\_}4)=\#C(\mathbb{F}_{\_}4)-2T_{\_}{A,4}-T_{\_}{E^{\prime},4}=\#C(\mathbb{F}_{\_}4)-14-T_{\_}{E^{\prime},4}\leq\#C(\mathbb{F}_{\_}4)-10;$
consequently, $C$ must be the unique curve with #$\#C(\mathbb{F}_{\_}4)=10$ and $J$ belongs to the isogeny class 2.4.f_n. However, in this case #$\#J(\mathbb{F}_{\_}4)=55$ is odd, and so by class field theory (as in [5, (7.5)]) $F$ cannot admit a nonconstant unramified quadratic extension, a contradiction.
7. $m=2$ and $q=3$
We prove Proposition 4.1 for $q=3$.
7.1. Analysis of the Prym
As in the case $q=4$, up to isogeny we can write
$A\sim E^{\dim(A)-\dim(B)}\times B,$
where $E$ is an elliptic curve with #$\#E(\mathbb{F}_{\_}3)=1$ in the isogeny class 1.3.ad and $B$ is an abelian variety not containing $E$ as a factor and satisfying #$\#B(\mathbb{F}_{\_}3)=2$. By Lemma 2.1, we have #$\#B(\mathbb{F}_{\_}3)\geq 1.359^{\dim(B)}$ and so $\dim(B)\leq 2$. By this LMFDB query, we obtain two possibilities for the isogeny class of $B$: an elliptic curve $B_{\_}1$ in 1.3.ac or a simple abelian surface $B_{\_}2$ in 2.3.ae_i. We have
$T_{\_}{E,3}=3,\qquad T_{\_}{B_{\_}1,3}=2,\qquad T_{\_}{B_{\_}2,3}=4,$
$T_{\_}{E,3^{2}}=3,\qquad T_{\_}{B_{\_}1,3^{2}}=-2,\qquad T_{\_}{B_{\_}2,3^{2}}=0,$
and therefore
(7.1) $(T_{\_}{A,3},T_{\_}{A,3^{2}})=\begin{cases}(3\dim(A)-1,3\dim(A)-5)&B=B_{\_}1,\\ (3\dim(A)-2,3\dim(A)-6)&B=B_{\_}2.\end{cases}$
7.2. Constant case
As in the case $q=4$, the fact that $53^{2}>$ 5 means that if $F^{\prime}/F$ is constant, we must have $d=2$.
When $A\sim E^{\dim(A)-\dim(B_{\_}1)}\times B_{\_}1$, by Lemma 2.3 we get
$3g+3=\#C(\mathbb{F}_{\_}3)\leq 1.91g+5.52,$
and we deduce that $g\leq 2$.
When $A\sim E^{\dim(A)-\dim(B_{\_}2)}\times B_{\_}2$, by Lemma 2.3 we get
$3g+2=\#C(\mathbb{F}_{\_}3)\leq 1.91g+5.52,$
and we deduce that $g\leq 3$. However, Table 1 indicates that for $g=3$, #$\#C(\mathbb{F}_{\_}3)\leq 10<11=3g+2$; we thus again obtain $g\leq 2$.
If $g=1$, then $A\sim E$ and so $J$ must belong to the isogeny class 1.3.d. If $g=2$, then $A$ cannot be isogenous to $E\times B_{\_}1$, as then its quadratic twist would belong to the isogeny class 2.3.fm which does not contain a Jacobian; hence $A\sim B_{\_}2$ and $J$ belongs to 2.3.ei.
7.3. Purely geometric case
7.3.1. $g=0$
In this case we have that #$\#C(\mathbb{F}_{\_}3)=4$, so when $B\sim B_{\_}1$, by using Equation 7.1 and Lemma 2.8, we have
$43(g^{\prime}-g)-1=3\dim A-1=T_{\_}{A,3}\leq\#C(\mathbb{F}_{\_}3)=$ 4 .
Therefore, $g^{\prime}=1$ and $C^{\prime}$ is isogenous to $B_{\_}1$ itself.
In the case $B\sim B_{\_}2$,
$3(g^{\prime}-g)-2=3\dim A-2=T_{\_}{A,3}\leq\#C(\mathbb{F}_{\_}3)\leq 4$
Therefore, $g^{\prime}=2$ and $C^{\prime}$ is isogenous to $B_{\_}2$ itself.
7.3.2. $g=1$
For $g=1$, #$\#C(\mathbb{F}_{\_}3)\leq 7$, so in the first case ($B\sim B_{\_}1$) we have
$3(g^{\prime}-g)-2\leq 7$
from what we get $g^{\prime}\leq 3$.
  • For $g^{\prime}=2$, we have $J^{\prime}\sim J\times B$; according to this LMFDB query, the possible isogeny classes for $J^{\prime}$ are 2.3.ae_k, 2.3.ac_g, 2.3.ab_e, 2.3.a_c, 2.3.b_a.
  • For $g^{\prime}=3$, we have $J^{\prime}\sim J\times E\times B$ and according to this LMFDB query, the possible isogeny clases are 3.3.ae_k_as, 3.3.ad_f_ag, 3.3.ac_a_g.
In the second case ($B\sim B_{\_}2$), we have $g^{\prime}\leq 4$.
  • For $g^{\prime}=3$, we have $J^{\prime}\sim J\times B$; according to this LMFDB query, the possible isogeny classes for $J^{\prime}$ are 3.3.ae_l_ay, 3.3.ad_h_aq, 3.3.ac_d_ai, 3.3.ab_ab_a.
  • For $g^{\prime}=4$, we have $J^{\prime}\sim J\times E\times B$ and according to this LMFDB query, the only possible isogeny class is 4.3.ae_f_a_ag.
This confirms the values listed in Table 3.
7.3.3. $g\geq 2$ and $d=2$
By Equation 2.10,
$1.91g+5.52\geq\#C(\mathbb{F}_{\_}3)\geq T_{\_}{A,3}\geq 3\dim(A)-2\geq 3g-5,$
and so $g\leq 9$. By Lemma 2.3,
#$\#C(\mathbb{F}_{\_}{3^{2}})$ $\geq 2\#C(\mathbb{F}_{\_}3)+T_{\_}{A,3^{2}}-t\geq 2\#C(\mathbb{F}_{\_}3)+3\dim(A)-6-t\geq 2\#C(\mathbb{F}_{\_}3)+3g-9,$
$\geq 2T_{\_}{A,3}+3g-9=6\dim(A)-4+3g-9\geq 9g-19.$
From Table 1, we see that for $g=9$, #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 50<62=9g-19$; for $g=8$, #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 46<53=9g-19$; and for $g=7$, #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 43<44=9g-19$. We deduce that $g\leq 6$.
For $g=6$, we have $214\geq\#C(\mathbb{F}_{\_}3)\geq 3(g^{\prime}-g)-$ 2, so $g^{\prime}=11$ and $F^{\prime}/F$ is unramified. We may rule out this case as follows. Since #$\#C(\mathbb{F}_{\_}3)\geq 13$, $938\geq\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+$ 9, and #$\#C(\mathbb{F}_{\_}{3^{2}})\equiv\#C(\mathbb{F}_{\_}3)\pmod{2}$, the pair (#$(\#C(\mathbb{F}_{\_}3),\#C(\mathbb{F}_{\_}{3^{2}}))$ must be one of
[…]
The indicated LMFDB searches show that none of these cases can occur.
For $g=5$, we have #$\#C(\mathbb{F}_{\_}3)\geq 3g-5=10$ and $213\geq\#C(\mathbb{F}_{\_}3)\geq 3(g^{\prime}-g)-$ 2; the latter yields $g^{\prime}\leq 10$. We rule out the case $g^{\prime}=10$ as follows. Since #$\#C(\mathbb{F}_{\_}3)\geq 10$, $835\geq\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+$ 8, and #$\#C(\mathbb{F}_{\_}{3^{2}})\equiv\#C(\mathbb{F}_{\_}3)\pmod{2}$, the pair (#$(\#C(\mathbb{F}_{\_}3),\#C(\mathbb{F}_{\_}{3^{2}}))$ must be one of
[…]
The indicated LMFDB searches show that none of these cases can occur.
For $g=5$, it still remains to rule out $g^{\prime}=9$ as follows. Since #$\#C(\mathbb{F}_{\_}3)\geq 10$, $635\geq\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+$ 6, #$\#C(\mathbb{F}_{\_}{3^{2}})\equiv\#C(\mathbb{F}_{\_}3)\pmod{2}$, and we have already excluded some options for the pair (#$(\#C(\mathbb{F}_{\_}3),\#C(\mathbb{F}_{\_}{3^{2}}))$, this pair must be one of
[…]
The indicated LMFDB searches show that none of these cases can occur except possibly the first one, for which we have the candidate isogeny classes 5.3.g_ba_da_hh_nw and 5.3.g_ba_db_hh_ob for $J$. The second one has the property that #$\#J(\mathbb{F}_{\_}3)$ is odd, so a curve with Jacobian $J$ cannot admit an étale double cover. Using the following lemma, we deduce that the case $g=5$ cannot occur.
Lemma 7.1.
The isogeny class 5.3.g_ba_da_hh_nw does not contain the isogeny class of a Jacobian.
Proof.
(Suggested by Everett Howe) Suppose the contrary that the isogeny class contains the Jacobian of a curve $C$. The “resultant 2” criterion [2, Theorem 4.5] shows that $C$ must be a double cover of a curve whose Jacobian belongs to one of the isogeny classes 2.3.d_i or 3.3.d_j_s; however, the former does not contain a Jacobian (e.g., by the “resultant 1” criterion). The latter contains the Jacobian of a unique curve $C_{\_}0$; the double cover $C\to C_{\_}0$ must be étale. We may now run over étale double covers of $C_{\_}0$ to see that none of them have the correct number of $\mathbb{F}_{\_}3$-rational points; this computation can be found in the notebook q3geometric.ipynb listed in Table 4. ∎
For $g=4$, Table 1 gives $212\geq\#C(\mathbb{F}_{\_}3)\geq 3(g^{\prime}-g)-$ 2, so $g^{\prime}-g\leq 4$ and $g^{\prime}\leq 8$. For the exhaust, we record the constraint
(7.2) $(q,g,g^{\prime})=(3,4,7)\Longrightarrow\#C(\mathbb{F}_{\_}4)\geq 3,\#C(\mathbb{F}_{\_}{3^{2}})\geq 17.$
We rule out $g^{\prime}=8$ as follows: we would have $1012\geq\#C(\mathbb{F}_{\_}3)\geq 3\cdot 4-2=$ 10 and $30\geq\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+3\cdot 4-6-2\geq 24$. The pair (#$(\#C(\mathbb{F}_{\_}3),\#C(\mathbb{F}_{\_}{3^{2}}))$ must be one of
[…]
The indicated LMFDB searches show that none of these cases can occur.
For $g=3$, Table 1 gives $210\geq\#C(\mathbb{F}_{\_}3)\geq 3(g^{\prime}-g)-$ 2, so $g^{\prime}-g\leq 4$ and $g^{\prime}\leq 7$. For the exhaust, we record the constraint
(7.3) $(q,g,g^{\prime})=(3,3,6)\Longrightarrow\#C(\mathbb{F}_{\_}3)\geq 7,\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}{3^{2}})+1.$
We rule out $g^{\prime}=7$ as follows: we would have $10\geq\#C(\mathbb{F}_{\_}3)\geq 10$ and #$\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+3(g^{\prime}-g^{\prime})-6-2(g^{\prime}-2g+1)=22$; but this LMFDB query shows that there are exactly two isomorphism classes of curves $C$ of genus 3 over $\mathbb{F}_{\_}3$ with #$\#C(\mathbb{F}_{\_}3)=10$, and both satisfy #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 12$.
For $g=2$, Table 1 gives $28\geq\#C(\mathbb{F}_{\_}3)\geq 3(g^{\prime}-g)-$ 2, so $g^{\prime}-g\leq 3$ and $g^{\prime}\leq 5$. We rule out $g^{\prime}=5$ as follows by noting that $8\geq\#C(\mathbb{F}_{\_}3)\geq 7$ and arguing as follows.
  • If #$\#C(\mathbb{F}_{\_}3)=8$, then #$\#C(\mathbb{F}_{\_}{3^{2}})\geq 15$. However, this LMFDB query shows that the possible isogeny classes for the Jacobian of a curve $C$ with #$\#C(\mathbb{F}_{\_}3)=8$ are 2.3.ei, 2.3.ej, 2.3.ek, and this shows that #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 14$.
  • If #$\#C(\mathbb{F}_{\_}3)=7$, then #$\#C^{\prime}(\mathbb{F}_{\_}3)=0$ and so the covering $C^{\prime}\to C$ cannot have any $\mathbb{F}_{\_}3$-rational ramification points. As in [5, (7.6), (7.7)], we must have $2\#C(\mathbb{F}_{\_}3)\leq\#C^{\prime}(\mathbb{F}_{\_}{3^{2}})=\#C(\mathbb{F}_{\_}{3^{2}})-T_{\_}{A,3^{2}}$ and so #$\#C(\mathbb{F}_{\_}{3^{2}})\geq 2\#C(\mathbb{F}_{\_}3)+T_{\_}{A,3^{2}}\geq 17$. However, this LMFDB query shows that the possible isogeny classes for the Jacobian of a curve $C$ with #$\#C(\mathbb{F}_{\_}3)=7$ are 2.3.df, 2.3.dg, 2.3.dh, and this shows that #$\#C(\mathbb{F}_{\_}{3^{2}})\leq 16$.
7.3.4. $g\geq 2$ and $d=3$
In the case that $B\sim B_{\_}1$, by the given conditions, Riemann–Hurwitz, Equation 7.1, and Lemma 2.3,
$6g-7\leq 3(g^{\prime}-g)-1=T_{\_}{A,3}\leq\#C(\mathbb{F}_{\_}3)\leq 1.91g+5.52,$
which implies $g=2,3$. For $g=3$, we have #$\#C(\mathbb{F}_{\_}3)\leq 10$ from Table 1, but this implies $g^{\prime}\leq 6$ which is impossible. For $g=2$, Table 1 indicates #$\#C(\mathbb{F}_{\_}3)\leq 8$, which implies that $4\leq g^{\prime}\leq 5$; note that here $g^{\prime}=5$ forces #$\#C(\mathbb{F}_{\_}3)=8$.
When $B\sim B_{\_}2$, by the given conditions, Riemann–Hurwitz, Equation 7.1, and Lemma 2.3,
$6g-8\leq 3(g^{\prime}-g)-2=T_{\_}{A,3}\leq\#C(\mathbb{F}_{\_}3)\leq 1.91g+5.52,$
which implies $g=2,3$. Again, we see that $4\leq g^{\prime}\leq 5$ for $g=2$, and $g^{\prime}=7$ for $g=3$; in this case, $g^{\prime}=5$ forces #$\#C(\mathbb{F}_{\_}3)\geq 7$.
By Lemma 2.9, $F^{\prime}/F$ cannot be cyclic. As in section 2.5, the quadratic resolvent of its Galois closure $F^{\prime\prime}/F$ is purely geometric. The Prym of the quadratic resolvent is an elliptic curve $E^{\prime}$ satisfying
$0\leq\#C^{\prime\prime}(\mathbb{F}_{\_}3)=\#C(\mathbb{F}_{\_}3)-2T_{\_}{A,3}-T_{\_}{E^{\prime},3}\leq\#C(\mathbb{F}_{\_}3)-2(3(g^{\prime}-g)-2)-T_{\_}{E^{\prime},3},$
and $|T_{\_}{E^{\prime},3}|\leq 3$. For $g=3$, this yields $10\geq\#C(\mathbb{F}_{\_}3)\geq 17$; for $g=2,g^{\prime}=5$, we get $8\geq\#C(\mathbb{F}_{\_}3)\geq 11$. Hence $g=2,g^{\prime}=4$.
We now have $5\leq 8+T_{\_}{E^{\prime},3}\leq\#C(\mathbb{F}_{\_}3)\leq 8$, so $T_{\_}{E^{\prime},3}\leq 0$. Also, $C^{\prime\prime\prime}$ admits a cyclic étale cubic cover which is not a base extension from $C$, so by class field theory, #$\#E^{\prime}(\mathbb{F}_{\_}3)\equiv 0\pmod{3}$. Together this yields $T_{\_}{E^{\prime},3}=-2$ and #$\#C(\mathbb{F}_{\_}3)\geq 2T_{\_}{A,3}-T_{\_}{E^{\prime},3}\geq 6$; using the following lemma, we deduce that #$\#C(\mathbb{F}_{\_}3)=6$ and $T_{\_}{A,3}=4$.
Lemma 7.2.
Let $C$ be a genus $2$ curve over $\mathbb{F}_{\_}3$ admitting an étale double cover whose Prym variety $E^{\prime}$ satisfies $|T_{\_}{E^{\prime},3}|=2$. Then $C$ appears in Table 9; in particular, #$\#C(\mathbb{F}_{\_}3)\leq 6$.
Proof.
If $T_{\_}{E^{\prime},3}=2$, the the given double cover has relative quadratic twist 2. If $T_{\_}{E^{\prime},3}=-2$, then the quadratic twist of $E^{\prime}$ is the Prym variety of the relative quadratic twist of the cover, so the latter has relative class number 2. ∎
Since $C$ admits an étale cover, #$\#J(\mathbb{F}_{\_}3)$ is even. Since we also have #$\#C(\mathbb{F}_{\_}3)=6$, we may use this LMFDB query to see that $J$ belongs to one of 2.3.cc, 2.3.ce, 2.3.cg; this restricts $C$ to a list of 6 possibilities. We enumerate towers of coverings $C^{\prime\prime}\to C^{\prime\prime\prime}\to C$ where $C^{\prime\prime\prime}\to C$ is étale of degree 2 with Prym isogenous to $E^{\prime}$ (per Lemma 7.2 these can be found in Table 9) and $C^{\prime\prime}\to C^{\prime\prime\prime}$ is cyclic étale of degree 3, looking for cases where the latter has relative class number 4; the code for this can be found in the notebook q3geometric.ipynb listed in Table 4. We find exactly one instance, which does give rise to an extension as described in the following lemma.
Lemma 7.3.
There exists a noncyclic cubic extension $F^{\prime}/F$ of function fields over $\mathbb{F}_{\_}3$ with relative class number $2$, corresponding to a map between the curves
$C\colon$ $y^{2}=x^{6}+x^{5}+2x^{4}+2x^{2}+2x+1,$
$C^{\prime}\colon$ $y^{2}=x^{9}+x^{8}+x^{7}+2x^{3}+x^{2}+2x.$
Proof.
The Jacobian $J$ of $C$ belongs to 2.3.cc. In particular, #$\#J(\mathbb{F}_{\_}3)$ is not divisible by 3, so $C$ does not admit an étale cyclic triple cover.
From the computation described above, we obtain an étale double cover $C^{\prime\prime\prime}\to C$ with Prym isogenous to $E^{\prime}$ and a unique étale cyclic triple cover $C^{\prime\prime}\to C^{\prime\prime\prime}$ with Prym isogenous to $B_{\_}2^{2}$. This uniqueness ensures that $C^{\prime\prime}\to C$ is Galois; this cover cannot be abelian because $C$ admits no étale cyclic triple cover, so its Galois group is $S_{\_}3$. Set $C^{\prime}:=C^{\prime\prime}/S_{\_}2$; then $C^{\prime}\to C$ is a triple cover with Prym isogenous to $B_{\_}2$. In particular, the Jacobian $J^{\prime}$ of $C^{\prime}$ belongs to 4.3.acccao, which contains a unique Jacobian; this confirms the equation for $C^{\prime}$.
For completeness, we also exhibit explicit equations for the map $C^{\prime}\to C$ (provided by Everett Howe):
$x$ $\mapsto\frac{x^{3}+2x+2}{x^{3}+2x^{2}+1},$
$y$ $\mapsto\frac{x^{4}+x^{3}+x^{2}+2x+1}{x^{9}+2x^{6}+1}y.\qed$
7.3.5. $g\geq 2$ and $d>3$
As in section 7.3.4, we have $g=2$ and
$7\leq 3(d-1)-2\leq 3(g^{\prime}-g)-2\leq T_{\_}{A,3}\leq\#C(\mathbb{F}_{\_}3)\leq 8$
It follows that $d=4$, $g=2$, $g^{\prime}=5$, and $F^{\prime}/F$ is unramified; moreover, #$\#C(\mathbb{F}_{\_}3)\in\{7,8\}$, which restricts $J$ to one of six isogeny classes listed in section 7.3.3. We need to show that no extensions occur in this case.
We first rule out the possibility that $F^{\prime}/F$ admits an intermediate subfield $F^{\prime\prime}$; note that this includes the case where $F^{\prime}/F$ is cyclic. In this case, the extension $F^{\prime\prime}/F$ would have relative class number 1 or 2; however, the former is impossible because [5, Table 4] implies #$\#C(\mathbb{F}_{\_}3)\in\{3,5\}$, while the latter is impossible since Lemma 7.2 implies #$\#C(\mathbb{F}_{\_}3)\leq 6$.
We assume hereafter that $F^{\prime}/F$ has no intermediate subfield, and set notation as in section 2.5; in particular, $F^{\prime\prime}/F$ is now the Galois closure of $F^{\prime}/F$. As in [6, Lemma 4.5], we can use the point counts of $C$ and $C^{\prime}$ to see that the Galois group of $F^{\prime\prime}/F$ must contain a 4-cycle, and so must equal $S_{\_}4$.
If the quadratic resolvent $F^{\prime\prime\prime}/F$ is constant, then $C_{\_}{\mathbb{F}_{\_}9}$ admits a cyclic cubic étale cover which is not a base extension from $C$, so as in [6, Remark 5.2] we have #$\#J(\mathbb{F}_{\_}9)/\#J(\mathbb{F}_{\_}3)\equiv 0\pmod{3}$; this excludes the isogeny classes 2.3.dg, 2.3.dh, 2.3.ei, 2.3.ek. We handle the remaining cases using splitting sequences as in [6].
  • If $J$ is in 2.3.df, then
    (#$(\#C(\mathbb{F}_{\_}{3^{i}}))_{\_}{i=1}^{2}=(7,11),\qquad(\#C^{\prime}(\mathbb{F}_{\_}{3^{i}}))_{\_}{i=1}^{4}=(0,8,33,40).$
    Because $F^{\prime\prime\prime}/F$ is constant, the splitting type of a place of $F$ of odd (resp. even) degree must correspond to the cycle type of an odd (resp. even) permutation in $S_{\_}4$. By looking at places of degree 1 and 2 in $F$ and $F^{\prime}$, we see that the splitting sequence must begin $\{1^{4}\times 7\},\{1^{4},2^{2}\}$; however, this creates too many places of degree 4.
  • If $J$ is in 2.3.ej, similar logic about splitting sequences rules out #$\#C^{\prime}(\mathbb{F}_{\_}3)=1$, so we have
    (#$(\#C(\mathbb{F}_{\_}{3^{i}}))_{\_}{i=1}^{3}=(8,12,20),\qquad(\#C^{\prime}(\mathbb{F}_{\_}{3^{i}}))_{\_}{i=1}^{3}=(0,8,30).$
    This time, the splitting sequence begins $\{4\times 8\}$, so the 10 degree-3 places of $F^{\prime}$ sit over the 4 degree-3 places of $F$; however, the latter all have splitting types $4$ or $2+1^{2}$ which leaves room for at most 8 degree-3 places of $F^{\prime}$.
If the quadratic resolvent is purely geometric, then it corresponds to an étale double cover $C^{\prime\prime\prime}\to C$. Let $E^{\prime}$ be the Prym of this cover; again as in [6, Remark 5.2], the existence of an étale cyclic triple cover of $C^{\prime\prime\prime}$ not induced from $C$ ensures that #$\#E^{\prime}(\mathbb{F}_{\_}3)\equiv 0\pmod{3}$. By Lemma 7.2 we cannot have $T_{\_}{E^{\prime},3}=-2$; this leaves only the option $T_{\_}{E^{\prime},3}=1$. Since $C^{\prime\prime\prime}\to C$ is a Galois cover, #$\#C^{\prime\prime\prime}(\mathbb{F}_{\_}3)=\#C(\mathbb{F}_{\_}3)-T_{\_}{E^{\prime},3}$ must be even, forcing #$\#C(\mathbb{F}_{\_}3)=7$. By class field theory, #$\#J(\mathbb{F}_{\_}3)$ must be even; this restricts $J$ to the single isogeny class 2.3.dg, so
#$\#C(\mathbb{F}_{\_}3)=7,\qquad\#C^{\prime\prime\prime}(\mathbb{F}_{\_}3)=6,\qquad(\#C^{\prime}(\mathbb{F}_{\_}{3^{i}}))_{\_}{i=1}^{2}=(0,10).$
The splitting sequence begins $\{4\times 2,2^{2}\times 3\}$, but this forces the contradiction #$\#C^{\prime}(\mathbb{F}_{\_}{3^{2}})\geq 12$.
8. Explicit calculations for geometric covers
In light of Proposition 4.1, the explicit resolution of the relative class number 2 problem for function fields with $q>2$ is reduced to an exhaust over quadratic covers for the particular values of $q,g,g^{\prime}$ indicated in Table 2 for which $g\geq 2$. We perform this exhaust using explicit class field theory as implemented in Magma, as in [5]; we report the result in a series of tables listed in Table 2. In some cases, we use constraints on #$\#C(\mathbb{F}_{\_}q)$ and #$\#C(\mathbb{F}_{\_}{q^{2}})$ derived in our earlier analysis; these are also cited in Table 2.
This exhaust requires as input a list of systems of polynomial equations defining isomorphism class representatives for the curves of genus $g$ over $\mathbb{F}_{\_}q$. As of this writing, this data is available in LMFDB for all pairs $(q,d)$ thtat we require; however, some notes on sourcing are in order.
  • For $q=5,g=2$, all of the curves are hyperelliptic, so the algorithm of Howe [3] can be used to enumerate isomorphism class representatives.
  • For $q=3,g=4$, a census of curves was made by the authors of [1], who generously shared their data with LMFDB.
$q$ $g$ $g^{\prime}$ Bound Printed table File Code
5 2 3 - Table 5 tables/tableq5g2d2gg3.txt q5geometric.ipynb
4 2 3 - Table 6 tables/tableq4g2d2gg3.txt q4geometric.ipynb
4 2 4 (6.4) Table 7 tables/tableq4g2d2gg4.txt
4 3 5 (6.3) Table 8 tables/tableq4g3d2gg5.txt
3 2 3 - Table 9 tables/tableq3g2d2gg3.txt q3geometric.ipynb
3 2 4 - Table 10 tables/tableq3g2d2gg4.txt
3 3 5 - Table 11 tables/tableq3g3d2gg5.txt
3 3 6 (7.3) - -
3 4 7 (7.2) Table 12 tables/tableq3g4d2gg7.txt
Table 4. Tuples $(q,g,g^{\prime})$ for which we explicitly tabulate geometric extensions, with links to the corresponding tables.
$J$ $F$
2.5.ce $y^{2}+4x^{6}+2x^{5}+2x^{4}+x^{3}+2x^{2}+4x+3$
2.5.ace $y^{2}+3x^{6}+4x^{5}+4x^{4}+2x^{3}+4x^{2}+3x+1$
2.5.ci $y^{2}+4x^{6}+x^{5}+4x^{4}+4x^{2}+3x+3$
2.5.aci $y^{2}+3x^{6}+2x^{5}+3x^{4}+3x^{2}+x+1$
2.5.ack $y^{2}+4x^{6}+x^{5}+3x^{3}+x+4$
2.5.ck $y^{2}+3x^{6}+2x^{5}+x^{3}+2x+3$
2.5.ae $y^{2}+4x^{5}+x^{4}+x^{3}+x^{2}+x+4$
2.5.ae $y^{2}+3x^{5}+2x^{4}+2x^{3}+2x^{2}+2x+3$
2.5.acc $y^{2}+4x^{5}+3x^{3}+3x+3$
2.5.cc $y^{2}+3x^{5}+x^{3}+x+1$
2.5.cg $y^{2}+4x^{5}+3x^{4}+2x^{3}+3x^{2}+2x+1$
2.5.acg $y^{2}+3x^{5}+x^{4}+4x^{3}+x^{2}+4x+2$
2.5.ack $y^{2}+4x^{5}+2x^{4}+4x^{3}+2x^{2}+4x$
2.5.ck $y^{2}+3x^{5}+4x^{4}+3x^{3}+4x^{2}+3x$
2.5.aa $y^{2}+4x^{6}+2x^{5}+3x+1$
2.5.acg $y^{2}+4x^{5}+4x^{4}+x^{3}+2x^{2}+2x$
2.5.cg $y^{2}+3x^{5}+3x^{4}+2x^{3}+4x^{2}+4x$
2.5.ac $y^{2}+4x^{5}+4x^{4}+2x^{2}+4x$
2.5.ac $y^{2}+4x^{5}+4x^{4}+x^{3}+3x^{2}+x$
2.5.ak $y^{2}+4x^{5}+4x$
2.5.ak $y^{2}+4x^{5}+4x$
2.5.ag $y^{2}+4x^{6}+3x^{4}+x^{2}+2$
2.5.ag $y^{2}+4x^{6}+3x^{4}+x^{2}+2$
2.5.cc $y^{2}+4x^{6}+4x^{5}+4x^{4}+x^{2}+x+1$
2.5.acc $y^{2}+3x^{6}+3x^{5}+3x^{4}+2x^{2}+2x+2$
2.5.aac $y^{2}+4x^{6}+4x^{5}+3x^{4}+x^{2}+x+2$
Table 5. Purely geometric covers of type $(q,g,d,g^{\prime})=(5,2,2,3)$.
$J$ $F$
2.4.acj $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+x^{3}+x^{2}+x$
2.4.acj $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+x^{3}+x^{2}+x$
2.4.acj $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+x^{3}+x^{2}+x$
2.4.cj $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+ax^{4}+x^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+x$
2.4.cj $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+ax^{4}+x^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+x$
2.4.cj $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+ax^{4}+x^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+x$
2.4.acd $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+ax^{5}+ax^{4}+ax^{2}+a$
2.4.acd $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a+1}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+ax^{4}+ax^{2}+1$
2.4.acb $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+ax$
2.4.aab $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+ax^{5}+ax^{3}+x^{2}+x$
2.4.aab $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{3}+x^{2}+x$
2.4.ab $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+ax^{5}+ax^{3}+x^{2}+a$
2.4.ab $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a+1}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{3}+x^{2}+a+1$
2.4.ad $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{3}+ax^{2}+ax$
2.4.ad $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{3}+ax^{2}+ax$
2.4.ad $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+ax^{5}+ax^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x$
2.4.ad $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+ax^{5}+ax^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x$
2.4.ah $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+x^{5}+x^{3}+x^{2}+ax+a$
2.4.ah $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+x^{5}+ax^{4}+x^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+ax+a+1$
2.4.acf $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a+1}\right)y+x^{5}+ax^{4}+ax^{3}+ax^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x+1$
2.4.acf $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+x^{5}+ax^{4}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{3}+ax^{2}+ax+a$
2.4.ach $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a+1}\right)y+ax^{5}+x^{3}+ax+a$
2.4.ach $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+x^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x+a+1$
2.4.cb $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x}\right)y+x^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{4}+ax$
2.4.cd $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+ax^{5}+a+1$
2.4.cd $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a+1}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+a$
2.4.cf $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a+1}\right)y+x^{5}+ax^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x+a$
2.4.cf $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+x^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{3}+ax+a+1$
2.4.ch $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a+1}\right)y+ax^{5}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{4}+x^{3}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+ax+a+1$
2.4.ch $y^{2}+\mathopen{}\mathclose{{}\left(x^{2}+x+a}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{5}+ax^{4}+x^{3}+ax^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x+a$
Table 6. Purely geometric covers of type $(q,g,d,g^{\prime})=(4,2,2,3)$.
$J$ $F$
2.4.di $y^{2}+xy+x^{5}+x^{3}+x$
Table 7. Purely geometric covers of type $(q,g,d,g^{\prime})=(4,2,2,4)$.
$J$ $F$
3.4.cfe $y^{2}+\mathopen{}\mathclose{{}\left(ax^{3}+ax^{2}}\right)y+x^{7}+x^{6}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x$
3.4.cfu $y^{2}+\mathopen{}\mathclose{{}\left(ax^{3}+ax^{2}}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{7}+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x$
3.4.chm $y^{2}+\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(a+1}\right)x^{4}+ax^{3}+ax^{2}}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x+a$
3.4.chm $y^{2}+\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(a+1}\right)x^{4}+ax^{3}+ax^{2}}\right)y+\mathopen{}\mathclose{{}\left(a+1}\right)x^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x$
3.4.enbc $y^{2}+\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(a+1}\right)x^{4}+ax^{3}+ax^{2}}\right)y+x^{3}+ax^{2}+\mathopen{}\mathclose{{}\left(a+1}\right)x+a$
Table 8. Purely geometric covers of type $(q,g,d,g^{\prime})=(4,3,2,5)$.
$J$ $F$
2.3.aac $y^{2}+2x^{5}+x$
2.3.aac $y^{2}+2x^{5}+x$
2.3.aac $y^{2}+2x^{5}+x$
2.3.aac $y^{2}+2x^{5}+x$
2.3.aa $y^{2}+2x^{5}+1$
2.3.aa $y^{2}+x^{5}+2$
2.3.ac $y^{2}+2x^{5}+2x$
2.3.ac $y^{2}+2x^{5}+2x$
2.3.ae $y^{2}+2x^{6}+x^{5}+2x+1$
2.3.acc $y^{2}+x^{5}+2x^{4}+2x^{3}+2x^{2}+2x+1$
2.3.acc $y^{2}+x^{6}+x^{5}+2x^{4}+2x^{2}+2x+1$
2.3.ace $y^{2}+x^{6}+2x^{5}+x^{4}+x+2$
2.3.acg $y^{2}+x^{5}+2x^{4}+x^{3}+2x^{2}+x$
2.3.acg $y^{2}+x^{6}+x^{5}+2x^{4}+2x^{3}+2x^{2}+x+1$
2.3.cc $y^{2}+2x^{5}+x^{4}+x^{3}+x^{2}+x+2$
2.3.cc $y^{2}+2x^{6}+2x^{5}+x^{4}+x^{2}+x+2$
2.3.ce $y^{2}+2x^{6}+x^{5}+2x^{4}+2x+1$
2.3.cg $y^{2}+2x^{5}+x^{4}+2x^{3}+x^{2}+2x$
2.3.cg $y^{2}+2x^{6}+2x^{5}+x^{4}+x^{3}+x^{2}+2x+2$
Table 9. Purely geometric covers of type $(q,g,d,g^{\prime})=(3,2,2,3)$.
$J$ $F$
2.3.bc $y^{2}+x^{6}+x^{5}+2x^{3}+x^{2}+x+2$
2.3.ce $y^{2}+2x^{6}+x^{5}+2x^{4}+2x+1$
2.3.ce $y^{2}+2x^{6}+x^{5}+2x^{4}+2x+1$
2.3.dg $y^{2}+2x^{6}+2x+2$
Table 10. Purely geometric covers of type $(q,g,d,g^{\prime})=(3,2,2,4)$.
$J$ $F$
3.3.aabg $xy^{3}+\mathopen{}\mathclose{{}\left(x+1}\right)y^{2}+y+x^{3}$
3.3.abae $y^{2}+2x^{7}+2x^{6}+2x^{3}+x^{2}+x+1$
3.3.abe $y^{2}+x^{7}+x^{6}+x^{3}+2x^{2}+2x+2$
3.3.abe $xy^{3}+\mathopen{}\mathclose{{}\left(2x^{2}+2}\right)y^{2}+\mathopen{}\mathclose{{}\left(x+1}\right)y+x^{4}$
3.3.ada $y^{2}+2x^{8}+2x^{7}+2x^{6}+x^{2}+2x+1$
3.3.ada $y^{2}+2x^{7}+2x^{5}+x^{4}+x^{3}+2x^{2}+1$
3.3.ada $y^{3}+x^{3}y+2x^{4}+x$
3.3.bbac $y^{2}+2x^{7}+x^{5}+2x^{4}+2x^{2}+2$
3.3.bbg $y^{2}+2x^{8}+x^{7}+2x^{6}+2x^{5}+x^{4}+2x^{3}+x^{2}+x$
3.3.bda $y^{2}+x^{8}+2x^{7}+x^{2}+x$
3.3.bde $y^{2}+2x^{8}+2x^{7}+2x^{6}+2x^{5}+x^{4}+x^{2}+x$
3.3.bdg $y^{2}+2x^{8}+2x^{6}+x^{5}+x^{3}+2x^{2}+2x+1$
3.3.cdc $y^{3}+xy^{2}+2x^{3}y+x^{4}+2x^{3}+x^{2}+x$
3.3.chi $y^{3}+\mathopen{}\mathclose{{}\left(2x^{2}+2x}\right)y^{2}+\mathopen{}\mathclose{{}\left(2x^{2}+x}\right)y+x^{4}+2x^{3}+x$
3.3.dho $y^{2}+2x^{8}+x^{7}+x^{6}+x^{5}+2x^{3}+2x^{2}+2x$
Table 11. Purely geometric covers of type $(q,g,d,g^{\prime})=(3,3,2,5)$.
$J$ $F$
4.3.dinba $y^{2}+x^{9}+2x^{8}+x^{5}+x^{2}+x+2$
4.3.ditbi $y^{2}+2x^{9}+x^{8}+x^{6}+x^{5}+x^{4}+2$
Table 12. Purely geometric covers of type $(q,g,d,g^{\prime})=(3,4,2,7)$.